Leon drove 270 miles to the lodge in the same time as Pat drove 330 miles to the lodge. If pat drove 10 miles per hour faster than Leon, then how fast did each of them drive?
Answer by stanbon(75887)

Leon drove 270 miles to the lodge in the same time as Pat drove 330 miles to the lodge. If pat drove 10 miles per hour faster than Leon, then how fast did each of them drive?
----------------
Leon DATA:
distance= 270 mi ; rate= x mph ; time=270/x hrs
--------
Pat DATA:
distance = 330 mi ; rate= x+10 mph ; time = 330/(x+10)
-----------
EQUATION:
time = time
270/x = 330/(x+10)
9(x+10) = 11x
9x+90 = 11x
2x=90
x=45 mph (Leon's rate)
x+10= 55 mph (Pat's rate)
============
Cheers,
Stan H.